# Quantum Simulation of Noisy Quantum Networks

### v1

@@ II.1 General remarks on quantum simulation of quantum networks
Quantum channel         CPM circuit (unitary $+$ noise)
@@
Noise and orchestration         CPM $+$ POVM circuits
@@ II.2.1 Quantum simulation of noise processes
Some works have analyzed the possibilities of quantum simulation using quantum devices in the context of open system dynamics, both in an analog [54, 55, 56, 57, 17] and digital [58, 59, 15, 60, 61, 62, 63, 64, 65] fashion. While the former approach is based on the emulation of quantum systems with other quantum systems such that the simulated system evolution is governed by a desired artificial Hamiltonian or Liovillian, the latter one makes use of other quantum systems to encode the desired information, and the simulation is carried out by appropriate quantum gates that invoke operations in a quantum computational way. Digital simulation is then closer to our work, where however, total or partial noise reduction and mitigation is used, and it requires dynamic control and high sampling costs.
@@
More broadly, the reproduction of quantum processes on quantum hardware has been considered in different contexts, fundamentally reducible to constructing quantum circuit analogues of operations—something already discussed in[47, 44]. In recent works [47, 44, 46, 77, 45, 78, 79], quantum simulation has been combined with variational optimization techniques (VQO) to find optimal solutions to specific network tasks or problems, where they also provide accessible and well-developed tools for simulating and variationally optimizing quantum network problems on classical or quantum computers [45]. These approaches also highlight the potential of quantum hardware to aid in quantum network development by leveraging adaptive learning strategies.
@@ III.1.1 Quantum channels, operations and representations
Any quantum operation that transforms density operators into density operators is called a quantum channel, i.e., ^$\hat{\mathcal{E}}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ where $\mathcal{L}(\mathcal{H})$ is the set of linear operators on $\mathcal{H}$ is a quantum channel if for all density operator $\rho$ then ^$\hat{\mathcal{E}}(\rho)$ is also a density operator. Note that any linear, trace-preserving and completely positive quantum map is a quantum channel. Usually, in particular in communication settings, a quantum channel is understood in a more narrow sense as a noise map that ideally corresponds to an identity operation. In the following, we do however not distinguish between quantum channels in this narrow sense and general noisy quantum operations (that ideally correspond to some unitary operation), as both are described by a completely positive map (CPM). In particular, all methods and results we describe in the following do apply to quantum channels and general CPMs.
@@
  • In the Kraus representation a quantum channel is characterized by the so-called Kraus operators, i.e., ^=$\hat{\mathcal{E}}=\{K_{i}\}$, which fulfill
@@
  • The Stinespring representation, which follows from the Stinespring (or dilation) theorem [96], states that any quantum channel can be understood as a unitary operation $\Lambda$ acting on a larger Hilbert space. A channel ^$\hat{\mathcal{E}}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ can be conceived as a unitary $\Lambda:\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}})\to\mathcal{L}(\mathcal{H}_{B}\otimes\mathcal{H}_{\bar{B}})$ that also maps the environment ($\bar{A}$) evolution, where part of the quantum state information is potentially leaked due system-environment interactions. The action of the map on an arbitrary state can be eventually described by tracing out $\mathcal{H}_{\bar{B}}$, i.e.,
    ^$\hat{\mathcal{E}}(\rho)=\text{tr}_{\bar{B}}\!\left[\Lambda\left(\left|0\right\rangle\!\left\langle 0\right|_{\bar{A}}\otimes\rho\right)\Lambda^{\dagger}\right].$ (3)
    The channel in the Stinespring representation is given by ^$\hat{\mathcal{E}}=(\Lambda,\left|0\right\rangle\!\left\langle 0\right|_{\bar{A}})$. In Appendix A we show how the Stinespring and Kraus representations uniquely relate.
@@
    $\Phi_{\hat{\mathcal{E}}}=\hat{\mathcal{E}}\otimes\mathbbm{1}\left(\left|\Phi\right\rangle\!\left\langle\Phi\right|\right),$ (4)
@@
    where ${p_{i}}$ forms a probability distribution and $\{\left|\psi_{i}\right\rangle\}$ represents a set of normalized states. By expressing $\left|\psi_{i}\right\rangle=\Omega_{i}\otimes\mathbbm{1}\left|\Phi\right\rangle$, where $\left\langle k\right|\Omega_{i}\left|l\right\rangle=\left\langle k,l|\psi_{i}\right\rangle$, it is evident that ${K_{i}=\sqrt{p_{i}},\Omega_{i}}$ constitutes a Kraus representation of $\hat{\mathcal{E}}$. Notably, this decomposition aligns with the non-uniqueness of the Kraus representation, as observed in Eq. (5). We denote the Kraus rank of $\hat{\mathcal{E}}$ as $r$, which is determined by the rank of $\Phi_{\hat{\mathcal{E}}}$, corresponding to the minimum number of elements in a Kraus representation of the channel $\hat{\mathcal{E}}$.
@@
    This representation is particularly significant as it facilitates channel concatenation through matrix multiplication. Specifically, the Liouville Superoperator of the concatenation of multiple channels corresponds to the matrix product of their respective Liouville Superoperators, i.e., if ^$\hat{\mathcal{E}}_{3}=\hat{\mathcal{E}}_{2}\circ\hat{\mathcal{E}}_{1}$, then $\mathcal{E}_{3}=\mathcal{E}_{2}\cdot\mathcal{E}_{1}$. Moreover, the action of a channel on an arbitrary state $\rho$ is given by $\mathcal{E}\left|\rho\right\rangle\!\rangle$, where $\left|\rho\right\rangle\!\rangle=\sum_{i,j}\left\langle i\right|\rho\left|j\right\rangle\left|ij\right\rangle$ represents the vectorized form of $\rho$.
@@
The distance between two channels, ^$\hat{\mathcal{E}}_{1}$ and ^$\hat{\mathcal{E}}_{2}$, can be evaluated using various metrics. In this work, we utilize the Choi fidelity, i.e.,
@@
The channel infidelity between two channels, ^$\hat{\mathcal{E}}_{1}$ and ^$\hat{\mathcal{E}}_{2}$, is defined as $1-F(\mathcal{E}_{1},\mathcal{E}_{2})$. Observe that, although experimentally unfeasible, this quantity can actually be accessed on a quantum computer, where an entangled state using an ancilla can be initially prepared, letting the ancilla untouched until the end. The noise introduced by the additional entangling gates and the additional ancilla required can be incorporated in the simulation to directly access the channel fidelity or the process matrix in a more efficient way, thereby reducing the overheads typically required, e.g., for full process tomography.
@@ III.2.2 Noisy channels
^$\hat{\mathcal{E}}(\rho)=p\rho+(1-p)\sigma_{z}\,\rho\,\sigma_{z},$ (8)
@@
^$\hat{\mathcal{E}}(\rho)=p\rho+\frac{1-p}{3}(\sigma_{z}\rho\,\sigma_{z}+\sigma_{x}\rho\,\sigma_{x}+\sigma_{y}\rho\,\sigma_{y}).$ (9)
Alternatively, it can be equivalently expressed as: ^$\hat{\mathcal{E}}(\rho)=p^{\prime}\rho+(1-p^{\prime})\rho\mathbbm{1}/2$, where $p^{\prime}=(4p-1)/3$. This transformation maps the state to a completely mixed state with probability $(1-p^{\prime})$, implying the loss of all state information with that probability. Notably, the depolarizing channel represents a comprehensive scenario for single-qubit channels, as any other channel or state can be transformed into a depolarizing form through random operations, a process known as depolarization.
@@ IV Tools for the simulation of channels, noisy operations and measurements
^$\hat{\mathcal{E}}(\rho)=\sum_{k}p_{k}\,\hat{\mathcal{E}}^{\text{ext}}_{k}(\rho),$ (11)
where $0<p_{k}\leq 1$, $\sum_{k}p_{k}=1$. Here, ^$\hat{\mathcal{E}}^{\text{ext}}_{k}$ denotes the extreme channels, which are those that cannot be expressed as convex combinations of other quantum channels [99, 100, 101]. Therefore, one can infer from Eq. (11) that $\hat{\mathcal{E}}$ can be realized by implementing ^$\hat{\mathcal{E}}^{\text{ext}}_{k}$ with probability $p_{k}$.
The Kraus rank of any extreme channel is at most $d$, where $d$ represents the dimension of the Hilbert space. Therefore an ancillary system of dimension $d$ is sufficient to implement any ^$\hat{\mathcal{E}}^{\text{ext}}$ and, thus, any quantum channel $\hat{\mathcal{E}}$. This insight leads to an overhead of $R_{n}=n$, meaning that the minimal size can be reduced to $N=2n$ qubits.
@@ IV.1.1 Relevant cases
[…]
Projective measurement channels.– Another class of channels that can be realized without ancillary systems are those that can be decomposed into a projective measurement followed by a correction operation. These channels possess a Kraus representation of the form ^=$\hat{\mathcal{M}}=\{U_{i}P_{i}\}$, where $P^{2}_{i}=P_{i}$ denotes a projector. Implementing channel $\hat{\mathcal{M}}$ involves performing a measurement described by $\{P_{i}\}$ followed by the correction operation $U_{i}$, eventually discarding the measurement outcome. An important example of this type is the quantum reset channel ($\hat{\mathcal{R}}$), which is a specific instance of an extreme channel that can be implemented without requiring ancillary systems, although this is not generally true for all extreme channels.
@@
where $\left|e\right\rangle$ is an orthogonal state to $\rho$. The erasure channel can be simulated with a single ancillary system in the pure state $\left|0\right\rangle_{a}$, by leaving the state untouched with probability $p$, and projecting the system to the state $\left|0\right\rangle\left|1\right\rangle_{a}$ with probability $(1-p)$, i.e.,
@@
Routine $\hat{\mathcal{P}}$.– Given a qudit channel ^=$\hat{\mathcal{E}}=\{K_{i}\}_{i=0}^{r-1}$, we initially apply a unitary gate $\Lambda$ that encodes $\{W^{\dagger}_{i}K_{i}\,\rho\,K_{i}^{\dagger}W_{i}\}_{i=0}^{r-1}$ in different orthogonal subspaces of the Hilbert space of the extended qudit system, where $W_{i}$ is a unitary gate. Subsequently, we perform a projective measurement $\{P_{i}\}_{i=0}^{r-1}$ into each of these subspaces. This measurement leaves the state of the entire system in the state $\frac{1}{p_{i}}W_{i}^{\dagger}K_{i}\rho K_{i}^{\dagger}W_{i}$ in the corresponding subspace with probability $p_{i}=\text{tr}[P_{i}\Lambda(\rho\oplus\boldsymbol{0})\Lambda^{\dagger}]$. We then implement a correction operation $\bar{W}_{i}$, which first returns the state of the qudit back into the data subspace and then inverts $W_{i}^{\dagger}$. At this stage, the state of the system is given by
@@ IV.3 POVM simulation
The approaches outlined above for simulating noisy channels and operations can also be extended to simulate noisy measurements, as described by a Positive Operator-Valued Measure (POVM). Given a POVM [47] $\mathcal{O}=\left\{O_{i}\geq 0\right\}_{i=0}^{p-1}$, it can always be implemented as a Von Neumann measurement acting on a larger Hilbert space. If the input state is considered to be given by $\rho\oplus\boldsymbol{0}$, then $\mathcal{O}$ can be performed by implementing channel ^=$\hat{\mathcal{E}}=\left\{\sqrt{O_{i}}\right\}_{i=0}^{p-1}$ as described previously. However, in this case, the outcome of the projective measurement on the qudit system is not erased but learned. Therefore, the dimension of the extended qudit system must satisfy $D\geq\sum_{i=0}^{p}\text{rank}(O_{i})$.

### v2

@@ II.1 General remarks on quantum simulation of quantum networks
Quantum channel      CPM circuit (unitary ++ noise)
@@
Noise and orchestration      CPM ++ POVM circuits
@@ II.2.1 Quantum simulation of noise processes
Some works have analyzed the possibilities of quantum simulation using quantum devices in the context of open system dynamics, both in an analog [54, 55, 56, 57, 17] and digital [58, 59, 15, 60, 61, 62, 63, 64, 65] fashion. While the former approach is based on the emulation of quantum systems with other quantum systems such that the simulated system evolution is governed by a desired artificial Hamiltonian or Liouvillian, the latter one makes use of other quantum systems to encode the desired information, and the simulation is carried out by appropriate quantum gates that invoke operations in a quantum computational way. Digital simulation is then closer to our work, where however, total or partial noise reduction and mitigation is used, and it requires dynamic control and high sampling costs.
@@
More broadly, the reproduction of quantum processes on quantum hardware has been considered in different contexts, fundamentally reducible to constructing quantum circuit analogues of operations—something already discussed in references [47, 44]. In recent works [47, 44, 46, 77, 45, 78, 79], quantum simulation has been combined with variational optimization techniques (VQO) to find optimal solutions to specific network tasks or problems, where they also provide accessible and well-developed tools for simulating and variationally optimizing quantum network problems on classical or quantum computers [45]. These approaches also highlight the potential of quantum hardware to aid in quantum network development by leveraging adaptive learning strategies.
@@ II.2.3 Comparison to network benchmarking
II.2.4 Comparison to quantum error mitigation approaches
Much effort has been focused on error or noise mitigation in quantum computing, aiming to recover the output of an ideal noiseless circuit despite hardware imperfections [96, 97, 98, 99, 100, 101, 102, 103]. These methods aim to suppress or extrapolate away unwanted device noise, typically at the cost of overhead in sampling or assumptions about the noise models.
Our work takes an opposite viewpoint: we aim not to remove hardware noise but to reshape, increase and exploit it so that the simulator reproduces a target noisy channel, such as a realistic memory or communication link in a quantum network, which can be expected be stronger than the inherent computer noise. The success criterion is therefore the fidelity with a desired noisy map, rather than the recovery of an ideal unitary. In many relevant cases (e.g., Pauli-type channels) the existing tools, like randomized compiling [101], can actually help to transform the native device noise into the desired one directly, helping homogenize the hardware noise and simplifying this transformation.
Thus our approach is complementary to standard mitigation: while mitigation tries to approach noiseless computation, we deliberately tailor imperfections to emulate the real-world noise of network components, turning a limitation of NISQ devices into a resource for quantum network simulation.
@@ III.1.1 Quantum channels, operations and representations
Any quantum operation that transforms density operators into density operators is called a quantum channel, i.e., $\hat{\mathcal{E}}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ where $\mathcal{L}(\mathcal{H})$ is the set of linear operators on $\mathcal{H}$ is a quantum channel if for all density operator $\rho$ then $\hat{\mathcal{E}}(\rho)$ is also a density operator. Note that any linear, trace-preserving and completely positive quantum map is a quantum channel. Usually, in particular in communication settings, a quantum channel is understood in a more narrow sense as a noise map that ideally corresponds to an identity operation. In the following, we do however not distinguish between quantum channels in this narrow sense and general noisy quantum operations (that ideally correspond to some unitary operation), as both are described by a completely positive map (CPM). In particular, all methods and results we describe in the following do apply to quantum channels and general CPMs.
@@
  • In the Kraus representation a quantum channel is characterized by the so-called Kraus operators, i.e., $\hat{\mathcal{E}}=\{K_{i}\}$, which fulfill
@@
  • The Stinespring representation, which follows from the Stinespring (or dilation) theorem [104], states that any quantum channel can be understood as a unitary operation $\Lambda$ acting on a larger Hilbert space. A channel $\hat{\mathcal{E}}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ can be conceived as a unitary $\Lambda:\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}})\to\mathcal{L}(\mathcal{H}_{B}\otimes\mathcal{H}_{\bar{B}})$ that also maps the environment ($\bar{A}$) evolution, where part of the quantum state information is potentially leaked due system-environment interactions. The action of the map on an arbitrary state can be eventually described by tracing out $\mathcal{H}_{\bar{B}}$, i.e.,
    $\hat{\mathcal{E}}(\rho)=\text{tr}_{\bar{B}}\!\left[\Lambda\left(\left|0\right\rangle\!\left\langle 0\right|_{\bar{A}}\otimes\rho\right)\Lambda^{\dagger}\right].$ (3)
    The channel in the Stinespring representation is given by $\hat{\mathcal{E}}=(\Lambda,\left|0\right\rangle\!\left\langle 0\right|_{\bar{A}})$. In Appendix A we show how the Stinespring and Kraus representations uniquely relate.
@@
    $\Phi_{\hat{\mathcal{E}}}=\mathbbm{1}\otimes\hat{\mathcal{E}}\left(\left|\Phi\right\rangle\!\left\langle\Phi\right|\right),$ (4)
@@
    where $p_{i}$} forms a probability distribution and $\{\left|\psi_{i}\right\rangle\}$ represents a set of normalized states. By expressing $\left|\psi_{i}\right\rangle=\mathbbm{1}\otimes\Omega_{i}\left|\Phi\right\rangle$, where $\left\langle k\right|\Omega_{i}\left|l\right\rangle=\left\langle k,l|\psi_{i}\right\rangle$, it is evident that $K_{i}=\sqrt{p_{i}},\Omega_{i}$} constitutes a Kraus representation of $\hat{\mathcal{E}}$. Notably, this decomposition aligns with the non-uniqueness of the Kraus representation, as observed in Eq. (5). We denote the Kraus rank of $\hat{\mathcal{E}}$ as $r$, which is determined by the rank of $\Phi_{\hat{\mathcal{E}}}$, corresponding to the minimum number of elements in a Kraus representation of the channel $\hat{\mathcal{E}}$.
@@
    This representation is particularly significant as it facilitates channel concatenation through matrix multiplication. Specifically, the Liouville Superoperator of the concatenation of multiple channels corresponds to the matrix product of their respective Liouville Superoperators, i.e., if $\hat{\mathcal{E}}_{3}=\hat{\mathcal{E}}_{2}\circ\hat{\mathcal{E}}_{1}$, then $\mathcal{E}_{3}=\mathcal{E}_{2}\cdot\mathcal{E}_{1}$. Moreover, the action of a channel on an arbitrary state $\rho$ is given by $\mathcal{E}\left|\rho\right\rangle\!\rangle$, where $\left|\rho\right\rangle\!\rangle=\sum_{i,j}\left\langle i\right|\rho\left|j\right\rangle\left|ij\right\rangle$ represents the vectorized form of $\rho$.
@@
The distance between two channels, $\hat{\mathcal{E}}_{1}$ and $\hat{\mathcal{E}}_{2}$, can be evaluated using various metrics. In this work, we utilize the Choi fidelity, i.e.,
@@
The channel infidelity between two channels, $\hat{\mathcal{E}}_{1}$ and $\hat{\mathcal{E}}_{2}$, is defined as $1-F(\mathcal{E}_{1},\mathcal{E}_{2})$.
Observe that, although experimentally unfeasible, this quantity can actually be accessed on a quantum computer, where an entangled state using an ancilla can be initially prepared, letting the ancilla untouched until the end. The noise introduced by the additional entangling gates and the additional ancilla required can be incorporated in the simulation to directly access the channel fidelity or the process matrix in a more efficient way, thereby reducing the overheads typically required, e.g., for full process tomography. Alternatively, process fidelity estimation techniques could be used in practice to test the quality of the simulations.
@@ III.2.2 Noisy channels
$\hat{\mathcal{E}}(\rho)=p\rho+(1-p)\sigma_{z}\,\rho\,\sigma_{z},$ (8)
@@
$\hat{\mathcal{E}}(\rho)=p\rho+\frac{1-p}{3}(\sigma_{z}\rho\,\sigma_{z}+\sigma_{x}\rho\,\sigma_{x}+\sigma_{y}\rho\,\sigma_{y}).$ (9)
Alternatively, it can be equivalently expressed as: $\hat{\mathcal{E}}(\rho)=p^{\prime}\rho+(1-p^{\prime})\rho\mathbbm{1}/2$, where $p^{\prime}=(4p-1)/3$. This transformation maps the state to a completely mixed state with probability $(1-p^{\prime})$, implying the loss of all state information with that probability. Notably, the depolarizing channel represents a comprehensive scenario for single-qubit channels, as any other channel or state can be transformed into a depolarizing form through random operations, a process known as depolarization.
@@ IV Tools for the simulation of channels, noisy operations and measurements
$\hat{\mathcal{E}}(\rho)=\sum_{k}p_{k}\,\hat{\mathcal{E}}^{\text{ext}}_{k}(\rho),$ (11)
where $0<p_{k}\leq 1$, $\sum_{k}p_{k}=1$. Here, $\hat{\mathcal{E}}^{\text{ext}}_{k}$ denotes the extreme channels, which are those that cannot be expressed as convex combinations of other quantum channels [107, 108, 109]. Therefore, one can infer from Eq. (11) that $\hat{\mathcal{E}}$ can be realized by implementing $\hat{\mathcal{E}}^{\text{ext}}_{k}$ with probability $p_{k}$.
The Kraus rank of any extreme channel is at most $d$, where $d$ represents the dimension of the Hilbert space. Therefore an ancillary system of dimension $d$ is sufficient to implement any $\hat{\mathcal{E}}^{\text{ext}}$ and, thus, any quantum channel $\hat{\mathcal{E}}$. This insight leads to an overhead of $R_{n}=n$, meaning that the minimal size can be reduced to $N=2n$ qubits.
@@ IV.1.1 Relevant cases
where $\{p_{k}\}$ form a probability distribution and $U_{k}$ represents a unitary gate. These channels can be implemented by applying gate $U_{i}$ with probability $p_{i}$, without the necessity for ancillary systems. Noteworthy noisy channels falling into this category include bit flip noise ($\hat{\mathcal{X}}=\{p\,\mathbbm{1},(1-p)X\}$), phase damping ($\hat{\mathcal{Z}}=\left\{\sqrt{p_{k}}\,e^{i\theta_{k}Z}\right\}_{k}$), and depolarizing noise, discussed in Sec. III.1.
Projective measurement channels.– Another class of channels that can be realized without ancillary systems are those that can be decomposed into a projective measurement followed by a correction operation. These channels possess a Kraus representation of the form $\hat{\mathcal{M}}=\{U_{i}P_{i}\}$, where $P^{2}_{i}=P_{i}$ denotes a projector. Implementing channel $\hat{\mathcal{M}}$ involves performing a measurement described by $\{P_{i}\}$ followed by the correction operation $U_{i}$, eventually discarding the measurement outcome. An important example of this type is the quantum reset channel ($\hat{\mathcal{R}}$), which is a specific instance of an extreme channel that can be implemented without requiring ancillary systems, although this is not generally true for all extreme channels.
@@
where $\left|e\right\rangle$ is an orthogonal state to $\rho$. The erasure channel can be simulated with a single ancillary system in the pure state $\left|0\right\rangle_{a}$, by leaving the state untouched with probability $p$, and projecting the system to the state $\left|0\right\rangle\left|1\right\rangle_{a}$ with probability (1p)(1-p), i.e.,
@@
Routine $\hat{\mathcal{P}}$.– Given a qudit channel $\hat{\mathcal{E}}=\{K_{i}\}_{i=0}^{r-1}$, we initially apply a unitary gate $\Lambda$ that encodes $\{W^{\dagger}_{i}K_{i}\,\rho\,K_{i}^{\dagger}W_{i}\}_{i=0}^{r-1}$ in different orthogonal subspaces of the Hilbert space of the extended qudit system, where $W_{i}$ is a unitary gate. Subsequently, we perform a projective measurement $\{P_{i}\}_{i=0}^{r-1}$ into each of these subspaces. This measurement leaves the state of the entire system in the state $\frac{1}{p_{i}}W_{i}^{\dagger}K_{i}\rho K_{i}^{\dagger}W_{i}$ in the corresponding subspace with probability $p_{i}=\text{tr}[P_{i}\Lambda(\rho\oplus\boldsymbol{0})\Lambda^{\dagger}]$. We then implement a correction operation $\bar{W}_{i}$, which first returns the state of the qudit back into the data subspace and then inverts $W_{i}^{\dagger}$. At this stage, the state of the system is given by
@@ IV.3 POVM simulation
The approaches outlined above for simulating noisy channels and operations can also be extended to simulate noisy measurements, as described by a Positive Operator-Valued Measure (POVM). Given a POVM [47] $\mathcal{O}=\left\{O_{i}\geq 0\right\}_{i=0}^{p-1}$, it can always be implemented as a Von Neumann measurement acting on a larger Hilbert space. If the input state is considered to be given by $\rho\oplus\boldsymbol{0}$, then $\mathcal{O}$ can be performed by implementing channel $\hat{\mathcal{E}}=\left\{\sqrt{O_{i}}\right\}_{i=0}^{p-1}$ as described previously. However, in this case, the outcome of the projective measurement on the qudit system is not erased but learned. Therefore, the dimension of the extended qudit system must satisfy $D\geq\sum_{i=0}^{p}\text{rank}(O_{i})$.